OpenStudy (anonymous):

Perform the indicated operation. When no operation is indicated, simplify the rational expression (need the step please)Thanks in advance (10m^4)/(9a^2 ) x 3a^2 b

OpenStudy (keenan):

combine and mulitply

OpenStudy (anonymous):

is it \[\frac{10m^4}{9a^2}\times 3a^2b\]?

OpenStudy (anonymous):

if so the answer is \[\frac{10m^4b}{3}\]